### v2

@@ 5 Numerical experiments
In this section, we conduct numerical experiments to evaluate the performance and validate the efficiency of our ripALM in Algorithm 1 for solving the QROT problem (4.1) and the BPDN problem (4.9). Specifically, we conduct experiments from the following three aspects:
  • In Section 5.1, we compare our ripALM with Snipal [24] and ciPALM [58], which represent the latest absolute-type and relative-type inexact pALMs, respectively, for solving the QROT problem on image data. These comparisons aim to illustrate how different error criteria with different tolerance parameters influence the practical numerical performance of the inexact pALM.
  • In Section 5.2, we compare our ripALM with Gurobi, Eckstein and Silva’s method [14] (denoted by ES-iALM) and a dual alternating direction method of multipliers (dADMM) presented in Appendix B, for solving large-scale QROT problems. Gurobi is one of the state-of-the-art commercial solvers, while ES-iALM is a relative-type inexact version of the vanilla ALM. Moreover, dADMM is a widely used first-order method for large-scale constrained convex optimization; see, for example, [6, 17]. In addition, it is worth noting that many other first-order methods can also be applied to the QROT problem, such as the Chambolle–Pock algorithm [8], the Condat–Vũ algorithm [9, 52], and the linearized ALM [53, 55]. However, our preliminary tests show that these methods often perform comparably to, or worse than, dADMM. For this reason, and for the sake of simplicity, we omit them from our comparisons and only select the dADMM in Appendix B as a representative first-order method.
  • In Section 5.3, we compare our ripALM with SPGL1 [51], Mosek, and the dADMM presented in Appendix E, for solving the BPDN problem. SPGL1 is a state-of-the-art solver specifically designed for BPDN, while Mosek is a highly efficient commercial solver that can be applied to solve the second-order cone programming reformulation of the BPDN problem; see Appendix F for more details.
@@
Initial points. For the experiments in Section 5.1, we simply initialize all inexact pALMs at the origin to amplify the impact of inexactness. For the experiments in Section 5.2 and Section 5.3, we will employ a warm-start strategy for more efficiency. Indeed, our numerical experience (see, e.g., [24, 26, 58]) have suggested that a reasonably good initial point would benefit the practical performance of such a “pALM + Ssn” algorithmic framework. Therefore, for the experiments in Section 5.2, we first employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1) to generate an initial point for our ripALM; see Appendix C for more details. Specifically, we terminate this iBPGM as long as it produces a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ such that $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-3}$, or its iteration number reaches 500. Moreover, we also apply this warm-start strategy to ES-iALM in Section 5.2 to ensure a fair comparison. For the experiments in Section 5.3, we use dADMM in Appendix E to generate an initial point. Specifically, we terminate dADMM as long as it produces a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ such that $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-3}$, or its iteration number reaches 100. Note that the time consumed during the above warm-starting phase is included in the total computational time for both ripALM and ES-iALM. In addition, the initial error variable $\bm{w}^{0}$ in ripALM and ES-iALM is always set to $\bm{0}$.
@@
5.1 Comparison with SNIPAL and ciPALM
@@
In this part of experiments, we compare our ripALM with Snipal [24] and ciPALM [58] for solving the QROT problem (4.1) to illustrate how different error criteria with different tolerance parameters influence the numerical performance of the inexact pALM. Note that both ripALM and ciPALM are of relative-type and only require a single tolerance parameter $\rho$, as shown in (2.5) and (2.8), respectively, while Snipal is of absolute-type and requires two summable tolerance parameter sequences $\{\varepsilon_{k}\}$ and $\{\delta_{k}\}$, as shown in (2.7). For simplicity, in our comparisons, we set $\varepsilon_{k}=\varepsilon_{0}/(k+1)^{p},\;\delta_{k}=\delta_{0}/(k+1)^{q}$ with $\varepsilon_{0}=\delta_{0}\in\big\{0.01,1\big\}$ and $p,q\in\big\{1.1,\,2.1,\,3.1\big\}$ (hence, there are 18 combinations in total) for Snipal. For ripALM and ciPALM, we consider $\rho\in\left\{0.01,0.05,0.1,0.3,0.5,0.7,0.9,0.99,0.999\right\}$ (hence, there are 9 choices).
@@
We use images from the ClassicImages and MicroscopyImages classes in the DOTmark collection [44], which serves as a benchmark dataset for the OT problem and its variants, to generate QROT instances. Note that the images in the ClassicImages (or MicroscopyImages) class consist of 10 different images, each with different resolutions of $32\times 32$, $64\times 64$, $128\times 128$ and $512\times 512$. Thus, for each resolution, we can pair any two different images within a class and compute the QROT problem, resulting in 90 QROT problems (45 from ClassicImages and 45 from MicroscopyImages). Moreover, the cost matrix $C$ is obtained by calculating the squared Euclidean distances between pixels.
@@
Tables 2 and 3 present the average results (over 90 instances) for $64\times 64$ resolution with the regularization parameter $\lambda\in\{1,\,0.1\}$. From the results, we see that the performance of all algorithms depends on the choices of tolerance parameters, and with proper tuning, their performances can be comparable. This is indeed expected because all the algorithms essentially employ the same “pALM + Ssn” framework and differ only in their error criteria for solving subproblems. Notably, since ripALM and ciPALM involve only a single tolerance parameter $\rho\in[0,1)$, they would be more user-friendly and easier to tune through a simple one-dimensional grid search. This supports the main motivation for developing a relative-type error criterion. Moreover, one can see that ripALM always outperforms ciPALM in terms of the total number of Ssn iterations and exhibits a more robust performance, particularly when a larger $\rho$ is chosen, leading to a noticeable reduction in computational time. This difference may be attributed to the extra correction step required by ciPALM to guarantee convergence, which tends to considerably increase the number of Ssn iterations [^4] One possible reason is that the variable obtained after the correction step may no longer serve as a suitable initial point for the Ssn method in the next iteration. needed to satisfy its error criterion (2.8), even though these extra iterations do not contribute to the progress of the pALM iterations. These observations further support the advantage of the vanilla inexact pALM framework advocated in this paper. As shown in Tables 2 and 3, our ripALM clearly exhibits greater robustness and efficiency. In addition, its performance remains relatively stable across different values of $\rho$, with slight advantage for $\rho\geq 0.5$.
@@
Table 2: Comparisons between ripALM, Snipal and ciPALM under different choices of tolerance parameters for solving the QROT problem (4.1), where $\lambda=1$ and the instances are generated using images with the resolution of $64\times 64$ in the ClassicImages and MicroscopyImages classes from the DOTmark collection. In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time in seconds.
Snipal
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
(1, 1.1, 1.1) 7.48e-07 17 (51) 8.90 (0.01, 1.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 1.1, 2.1) 7.48e-07 17 (58) 9.76 (0.01, 1.1, 2.1) 7.48e-07 17 (70) 11.42
(1, 1.1, 3.1) 7.48e-07 17 (62) 10.23 (0.01, 1.1, 3.1) 7.48e-07 17 (72) 11.65
(1, 2.1, 1.1) 7.48e-07 17 (51) 8.89 (0.01, 2.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 2.1, 2.1) 7.48e-07 17 (58) 9.74 (0.01, 2.1, 2.1) 7.48e-07 17 (70) 11.43
(1, 2.1, 3.1) 7.48e-07 17 (62) 10.21 (0.01, 2.1, 3.1) 7.48e-07 17 (72) 11.66
(1, 3.1, 1.1) 7.48e-07 17 (51) 8.89 (0.01, 3.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 3.1, 2.1) 7.48e-07 17 (58) 9.75 (0.01, 3.1, 2.1) 7.48e-07 17 (70) 11.41
(1, 3.1, 3.1) 7.48e-07 17 (62) 10.22 (0.01, 3.1, 3.1) 7.48e-07 17 (72) 11.65
ripALM ciPALM
$\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
0.999 7.48e-07 17 (42) 8.01 0.999 7.48e-07 17 (164) 71.95
0.99 7.48e-07 17 (42) 8.02 0.99 7.55e-07 17 (163) 71.53
0.9 7.48e-07 17 (42) 8.03 0.9 7.40e-07 17 (156) 67.33
0.7 7.48e-07 17 (42) 8.10 0.7 7.48e-07 17 (142) 58.87
0.5 7.48e-07 17 (43) 8.18 0.5 7.48e-07 17 (125) 48.21
0.3 7.48e-07 17 (44) 8.32 0.3 7.48e-07 17 (105) 36.16
0.1 7.48e-07 17 (47) 8.68 0.1 7.48e-07 17 (79) 20.00
0.05 7.48e-07 17 (49) 8.93 0.05 7.48e-07 17 (70) 14.73
0.01 7.48e-07 17 (55) 9.59 0.01 7.48e-07 17 (65) 11.30
Table 3: Same as Table 2 but for $\lambda=0.1$.
Snipal
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
(1, 1.1, 1.1) 6.19e-07 20 (92) 27.72 (0.01, 1.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 1.1, 2.1) 6.19e-07 20 (100) 28.60 (0.01, 1.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 1.1, 3.1) 6.19e-07 20 (104) 29.11 (0.01, 1.1, 3.1) 6.19e-07 20 (115) 30.56
(1, 2.1, 1.1) 6.19e-07 20 (92) 27.74 (0.01, 2.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 2.1, 2.1) 6.19e-07 20 (100) 28.63 (0.01, 2.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 2.1, 3.1) 6.19e-07 20 (104) 29.10 (0.01, 2.1, 3.1) 6.19e-07 20 (115) 30.56
(1, 3.1, 1.1) 6.19e-07 20 (92) 27.72 (0.01, 3.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 3.1, 2.1) 6.19e-07 20 (100) 28.60 (0.01, 3.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 3.1, 3.1) 6.19e-07 20 (104) 29.09 (0.01, 3.1, 3.1) 6.19e-07 20 (115) 30.57
ripALM ciPALM
$\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
0.999 6.20e-07 20 (81) 26.68 0.999 6.19e-07 20 (230) 111.70
0.99 6.20e-07 20 (81) 26.68 0.99 6.19e-07 20 (232) 113.24
0.9 6.20e-07 20 (82) 26.72 0.9 6.19e-07 20 (222) 105.97
0.7 6.20e-07 20 (82) 26.78 0.7 6.19e-07 20 (200) 91.75
0.5 6.20e-07 20 (83) 26.85 0.5 6.19e-07 20 (181) 79.50
0.3 6.19e-07 20 (85) 27.00 0.3 6.19e-07 20 (156) 63.09
0.1 6.19e-07 20 (88) 27.39 0.1 6.19e-07 20 (122) 41.27
0.05 6.19e-07 20 (90) 27.68 0.05 6.19e-07 20 (115) 36.27
0.01 6.19e-07 20 (96) 28.39 0.01 6.19e-07 20 (108) 31.14
5.2 Comparison with Gurobi, dADMM and ES-iALM
@@
In this part of experiments, we compare our ripALM with Gurobi (version 12.0.1), dADMM and ES-iALM for solving large-scale QROT problems. For our ripALM, we set $\rho=0.99$ based on the numerical observations from the previous section. For ES-iALM, we adopt the same settings for $\rho$, $\sigma_{k}$}, and the termination conditions as those used in ripALM. Moreover, we employ FISTA [3] to solve the ALM-subproblems, with the maximum number of total inner iterations set to 15000.555Note that since ES-iALM does not involve a proximal term, the Ssn method may not be applicable to its subproblems, as the nonsingularity of the generalized Jacobian cannot be guaranteed. Therefore, we employ FISTA as the subsolver in ES-iALM. For Gurobi, we use its default termination conditions and set the corresponding termination tolerances as $10^{-6}$, aligning it with our tolerance. For dADMM, we initialize the penalty parameter as $\sigma_{0}=0.01\|C\|_{F}^{-1}$, and dynamically adjust it based on the primal-dual residuals, following the approach described in [57, Section 5.1], to achieve superior empirical performance. Moreover, we will terminate dADMM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-6}$, or its number of iterations reaches 10000.
@@
We follow [59, Section 5.1] to generate a random instance. Specifically, we first generate two discrete probability distributions $\big\{(a_{i},\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big\}$ and $\big\{(b_{j},\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big\}$. Here, $\bm{a}:=(a_{1},\cdots,a_{m})^{\top}$ and $\bm{b}:=(b_{1},\cdots,b_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval $(0,1)$ and further normalized such that $\sum^{m}_{i}a_{i}=\sum^{n}_{j}b_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
@@
We then generate a set of random instances with $m=n\in\{1000,\,2000,\dots,10000\}$. For each $m$, we generate 10 instances with different random seeds, and present the average numerical performances of Gurobi, dADMM, ES-iALM, and ripALM in Table 4 and Figure 1, with $\lambda\in\{1,\,0.1\}$. We observe that both ripALM and Gurobi are able to solve the tested problems accurately, in the sense that the residual $\Delta_{\mathrm{res}}^{\rm qrot}$ falls below $10^{-6}$. In contrast, ES-iALM often fails to attain solutions of comparable accuracy, even after reaching its maximum iteration limit and incurring substantially higher computational costs. This is mainly due to the limited efficiency of the first-order subsolver used in ES-iALM. Moreover, dADMM performs even worse: despite consuming a similar amount of computational time as ES-iALM, it yields solutions of considerably lower quality. In addition, we have also observed that Gurobi can be rather time-consuming and memory-consuming for large-scale problems. As an example, for the case where $m=n=10000$, a large-scale QP containing $10^{8}$ nonnegative variables and 20000 equality constraints was solved. One can see that Gurobi is around $2\!\sim\!4$ times slower than our ripALM. In addition, Gurobi may lack robustness, especially for solving large-scale problems. Indeed, as observed from Figure 1, the computational times taken by Gurobi vary a lot among 10 randomly generated instances. In conclusion, our ripALM shows potential for greater efficiency and robustness in solving large-scale QROT problems.
@@
Table 4: Numerical results of Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1). In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$; “#” denotes the number of iterations (the total number of Ssn/FISTA iterations is given in the bracket); “time” denotes the computational time.

### v3

@@ 5 Numerical experiments
In this section, we conduct numerical experiments to evaluate the performance of our ripALM in Algorithm 1 for solving the QROT problem (4.1) and the BPDN problem (4.9). Our experiments are organized as follows:
  • Section 5.1 reports experiments on QROT. We first compare ripALM with Snipal [19] and ciPALM [48], which represent the latest absolute-type and relative-type inexact pALMs, respectively, for solving QROT problems on image data. These comparisons are intended to illustrate how different error criteria and tolerance parameters affect the numerical performance of the inexact pALM. We then compare ripALM with Gurobi, Eckstein and Silva’s method [9] (denoted by ES-iALM), and a dual alternating direction method of multipliers (dADMM) presented in Appendix B, for solving large-scale QROT problems. Here, Gurobi is one of the state-of-the-art commercial solvers, ES-iALM is a relative-type inexact version of the vanilla ALM, and dADMM is a widely used first-order method for large-scale constrained convex optimization; see, e.g, [6, 12].
  • Section 5.2 reports experiments on BPDN. Similar to the QROT experiments, we first compare ripALM with Snipal and ciPALM to examine how different error criteria and tolerance parameters influence the numerical performance. We then compare ripALM with SPGL1 [44], Mosek, and the dADMM presented in Appendix E, for solving the BPDN problem on the UCI dataset. Here, SPGL1 is a state-of-the-art solver specifically designed for BPDN, while Mosek is a highly efficient commercial solver that can be used to solve the second-order cone programming reformulation of the BPDN problem; see Appendix F for more details.
@@
Initial points. For comparisons among inexact pALMs, we initialize all methods at the origin in order to isolate the impact of inexactness. For comparisons with other representative methods in Section 5.1 and Section 5.2, we will employ a warm-start strategy to improve the efficiency of ripALM. Specifically, for the experiments in Section 5.1, we first employ an inexact Bregman proximal gradient method (iBPGM), with Sinkhorn’s algorithm as a subsolver, for approximately solving the QROT problem (4.1) to generate an initial point for ripALM; see Appendix C for more details. We terminate iBPGM once it produces a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-3}$, or once the number of iterations reaches 500. Moreover, we also apply this warm-start strategy to ES-iALM in Section 5.1 to ensure a fair comparison. For the experiments in Section 5.2, we use dADMM in Appendix E to generate an initial point. We terminate dADMM once it produces a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-3}$, or once the iteration number reaches 100. Note that the time consumed by the above warm-starting phase is included in the total computational time for both ripALM and ES-iALM. In addition, the initial error variable $\bm{w}^{0}$ in ripALM and ES-iALM is always set to $\bm{0}$.
@@
5.1 Experiments on QROT
@@
In this subsection, we first compare ripALM with Snipal [19] and ciPALM [48] for solving the QROT problem (4.1) to illustrate how different error criteria and tolerance parameters affect the numerical performance of the inexact pALM. Notably, both ripALM and ciPALM are of relative-type and only require a single tolerance parameter $\rho$ in $[0,1)$, as shown in (2.5) and (2.8), respectively, while Snipal is of absolute-type and requires two summable tolerance parameter sequences $\{\varepsilon_{k}\}\subseteq\mathbb{R}_{+}$ and $\{\delta_{k}\}\subseteq\mathbb{R}_{+}$, as shown in (2.7). For simplicity, in our comparisons, we set $\varepsilon_{k}=\varepsilon_{0}/(k+1)^{p},\;\delta_{k}=\delta_{0}/(k+1)^{q}$ with $\varepsilon_{0}=\delta_{0}\in\big\{1,0.1,0.01\big\}$ and $p,q\in\big\{1.1,\,2.1,\,3.1\big\}$ (hence, there are 27 combinations in total) for Snipal. For ripALM and ciPALM, we choose $\rho\in\left\{0.01,0.05,0.1,0.3,0.5,0.7,0.9,0.99,0.999\right\}$ (9 choices).
@@
We use images of resolution $64\times 64$ from the ClassicImages class in the DOTmark collection [39], which serves as a benchmark dataset for the OT problem and its variants, to generate QROT instances. Specifically, the ClassicImages class contains 10 different images, and each QROT instance is generated by randomly selecting two different images from this class. The cost matrix $C$ is constructed by calculating the squared Euclidean distances between pixels.
@@
Table 2 presents the average results over 10 instances for the regularization parameter $\lambda\in\{1,\,0.1\}$. Since all three algorithms attain a similar solution accuracy and their computational times are proportional to the total number of Ssn iterations, we report only the numbers of outer (pALM) and inner (Ssn) iterations, denoted by “# (#)”, for a clearer comparison. From the results, it is not surprising to see that the performance of all algorithms depends on the choice of tolerance parameters. In particular, ripALM and ciPALM involve only a single tolerance parameter $\rho$ in the smaller interval $[0,1)$, and are therefore easier to tune through a simple one-dimensional grid search. In contrast, Snipal requires a three-dimensional grid search over parameters chosen from a much larger range. This supports the practical motivation for developing a relative-type error criterion that avoids prescribing summable tolerance sequences and is more convenient to implement. Moreover, ripALM consistently outperforms ciPALM in terms of the total number of Ssn iterations and exhibits more robust performance, especially when a larger $\rho\geq 0.5$ is used. This difference may be attributed to the additional correction step required by ciPALM to guarantee convergence, which tends to considerably increase the number of Ssn iterations [^3] One possible reason is that the variable obtained after the correction step may no longer serve as a suitable initial point for the Ssn method in the next iteration. needed to satisfy its error criterion (2.8), even though these extra iterations do not contribute to the progress of the pALM iterations. These observations further support the advantage of the vanilla inexact pALM framework advocated in this paper.
@@
Table 2: Comparisons among Snipal, ciPALM, and ripALM under different choices of tolerance parameters for solving the QROT problem (4.1) with $\lambda\in\{1,0.1\}$, where the test instances are generated using images of resolution $64\times 64$ from the ClassicImages class in the DOTmark collection.
$\lambda=1$
Snipal ciPALM ripALM
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $\rho$ # (#) $\rho$ # (#)
(1, 1.1, 1.1) 17 (45) (0.1, 1.1, 1.1) 17 (53) (0.01, 1.1, 1.1) 17 (58) 0.999 17 (165) 0.999 17 (37)
(1, 1.1, 2.1) 17 (51) (0.1, 1.1, 2.1) 17 (57) (0.01, 1.1, 2.1) 17 (61) 0.99 17 (165) 0.99 17 (37)
(1, 1.1, 3.1) 17 (55) (0.1, 1.1, 3.1) 17 (59) (0.01, 1.1, 3.1) 17 (63) 0.9 17 (162) 0.9 17 (38)
(1, 2.1, 1.1) 17 (45) (0.1, 2.1, 1.1) 17 (53) (0.01, 2.1, 1.1) 17 (58) 0.7 17 (139) 0.7 17 (38)
(1, 2.1, 2.1) 17 (51) (0.1, 2.1, 2.1) 17 (57) (0.01, 2.1, 2.1) 17 (61) 0.5 17 (126) 0.5 17 (39)
(1, 2.1, 3.1) 17 (55) (0.1, 2.1, 3.1) 17 (59) (0.01, 2.1, 3.1) 17 (63) 0.3 17 (100) 0.3 17 (40)
(1, 3.1, 1.1) 17 (45) (0.1, 3.1, 1.1) 17 (53) (0.01, 3.1, 1.1) 17 (58) 0.1 17 (75) 0.1 17 (42)
(1, 3.1, 2.1) 17 (51) (0.1, 3.1, 2.1) 17 (57) (0.01, 3.1, 2.1) 17 (61) 0.05 17 (65) 0.05 17 (44)
(1, 3.1, 3.1) 17 (55) (0.1, 3.1, 3.1) 17 (59) (0.01, 3.1, 3.1) 17 (63) 0.01 17 (56) 0.01 17 (48)
$\lambda=0.1$
Snipal ciPALM ripALM
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $\rho$ # (#) $\rho$ # (#)
(1, 1.1, 1.1) 19 (86) (0.1, 1.1, 1.1) 19 (96) (0.01, 1.1, 1.1) 19 (100) 0.999 19 (236) 0.999 19 (77)
(1, 1.1, 2.1) 19 (93) (0.1, 1.1, 2.1) 19 (99) (0.01, 1.1, 2.1) 19 (103) 0.99 19 (247) 0.99 19 (77)
(1, 1.1, 3.1) 19 (96) (0.1, 1.1, 3.1) 19 (101) (0.01, 1.1, 3.1) 19 (106) 0.9 19 (235) 0.9 19 (77)
(1, 2.1, 1.1) 19 (86) (0.1, 2.1, 1.1) 19 (96) (0.01, 2.1, 1.1) 19 (100) 0.7 19 (213) 0.7 19 (78)
(1, 2.1, 2.1) 19 (93) (0.1, 2.1, 2.1) 19 (99) (0.01, 2.1, 2.1) 19 (103) 0.5 19 (178) 0.5 19 (78)
(1, 2.1, 3.1) 19 (96) (0.1, 2.1, 3.1) 19 (101) (0.01, 2.1, 3.1) 19 (106) 0.3 19 (150) 0.3 19 (80)
(1, 3.1, 1.1) 19 (86) (0.1, 3.1, 1.1) 19 (96) (0.01, 3.1, 1.1) 19 (100) 0.1 19 (112) 0.1 19 (82)
(1, 3.1, 2.1) 19 (93) (0.1, 3.1, 2.1) 19 (99) (0.01, 3.1, 2.1) 19 (103) 0.05 19 (109) 0.05 19 (84)
(1, 3.1, 3.1) 19 (96) (0.1, 3.1, 3.1) 19 (101) (0.01, 3.1, 3.1) 19 (106) 0.01 19 (100) 0.01 19 (89)
@@
To further evaluate the numerical performance of ripALM, we compare ripALM with Gurobi (version 12.0.3), dADMM, and ES-iALM for solving large-scale QROT problems. For ripALM, we set $\rho=0.99$ based on the numerical observations from Table 2. For ES-iALM, we adopt the same settings for $\rho$, $\sigma_{k}$}, and the termination conditions as those used in ripALM. Moreover, we employ FISTA [3] to solve the ALM-subproblems, with the maximum number of total inner iterations set to 15000.444Since ES-iALM does not involve a proximal term in its subproblem, the Ssn method may not be applicable, as the nonsingularity of the generalized Jacobian cannot be guaranteed. Therefore, we employ FISTA as the subsolver in ES-iALM. For Gurobi, we use its default termination conditions and set the termination tolerances as $10^{-6}$, aligning it with our tolerance. For dADMM, we initialize the penalty parameter as $\sigma_{0}=0.01\|C\|_{F}^{-1}$, and dynamically adjust it based on the primal-dual residuals, following the approach described in [47, Section 5.1], to achieve superior performance. Moreover, we will terminate dADMM when it returns a point $(\bm{u}^{k},\bm{v}^{k},X^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\bm{v}^{k},X^{k})<10^{-6}$, or its number of iterations reaches 10000.
@@
We follow [49, Section 5.1] to generate a random instance. First, we generate two discrete probability distributions $\big\{(\alpha_{i},\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big\}$ and $\big\{(\beta_{j},\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big\}$. Here, $\bm{\alpha}:=(\alpha_{1},\cdots,\alpha_{m})^{\top}$ and $\bm{\beta}:=(\beta_{1},\cdots,\beta_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval $(0,1)$ and further normalized such that $\sum^{m}_{i}\alpha_{i}=\sum^{n}_{j}\beta_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
@@
We then generate a set of random instances with $m=n\in\{1000,\,2000,\dots,10000\}$. For each $m$, we generate 10 instances with different random seeds, and present the average results of Gurobi, dADMM, ES-iALM, and ripALM in Table 3, with $\lambda\in\{1,\,0.1\}$. We observe that both ripALM and Gurobi are able to solve the tested problems accurately, in the sense that the residual $\Delta_{\mathrm{res}}^{\rm qrot}$ falls below $10^{-6}$. In contrast, ES-iALM often fails to attain solutions of comparable accuracy, even after reaching its maximum iteration limit and incurring substantially higher computational costs. This is mainly due to the limited efficiency of the first-order subsolver used in ES-iALM. Moreover, dADMM performs even worse: despite consuming a similar amount of computational time as ES-iALM, it yields solutions of considerably lower quality. In addition, Gurobi can also be time-consuming and memory-consuming when problem size is large. For example, when $m=n=10000$, the resulting QP contains $10^{8}$ nonnegative variables and 20000 equality constraints. In this case, Gurobi is about $1.5\!\sim\!3$ times slower than our ripALM.
@@
Table 3: Comparisons among Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1) with $\lambda\in\{1,0.1\}$. In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\bm{v}^{k},X^{k})$; “#” denotes the number of iterations (the total number of Ssn/FISTA iterations is given in the bracket); “time” denotes the computational time.